rxcklesskaisher:

Which of the following is the solution to the equation 3 over 4y = 24? A) y = 64 B) y = 18 C) y = 32 D) y = 72

snowflake0531:

i think that you wrote the thing wrong, it's not possible if it's 3/(4y)=24

rxcklesskaisher:

im thinking its B

rxcklesskaisher:

thats what the question says so ig the computer dumb

aevans182:

did you mean 3/4y=24

snowflake0531:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @rxcklesskaisher thats what the question says so ig the computer dumb \(\color{#0cbb34}{\text{End of Quote}}\) is that what it exactly word for word says?

rxcklesskaisher:

rxcklesskaisher:

rxcklesskaisher:

that

snowflake0531:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @rxcklesskaisher that \(\color{#0cbb34}{\text{End of Quote}}\) lol, you wrote it wrong you shoulda just screenshotted it lmao

rxcklesskaisher:

i did

rxcklesskaisher:

hahah so funny..

snowflake0531:

if it's 3y/4 = 24 first, multiply 4 on each side what does that give you?

rxcklesskaisher:

100

KpopAsianDramaFanYT:

Solve for y by simplifying both sides of the equation, then isolating the variable. y = 32

snowflake0531:

and btw, i hate buz lmao

jhonyy9:

\[\frac{ 3 }{ 4 }y = 24 \] in this way

snowflake0531:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @rxcklesskaisher 100 \(\color{#0cbb34}{\text{End of Quote}}\) no 24 *4 is not 100, that's 25 it's 96

snowflake0531:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 and btw, i hate buz lmao \(\color{#0cbb34}{\text{End of Quote}}\) *buzz

snowflake0531:

anways

snowflake0531:

so that would make 3y=96

KpopAsianDramaFanYT:

y = 32

rxcklesskaisher:

oml

KpopAsianDramaFanYT:

what oml?

rxcklesskaisher:

ok so the answer is 96?!?!?

KpopAsianDramaFanYT:

no 32!!!

rxcklesskaisher:

im so lost

snowflake0531:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @rxcklesskaisher ok so the answer is 96?!?!? \(\color{#0cbb34}{\text{End of Quote}}\) no 3y=96 divide 3 by each side

snowflake0531:

just divide 96 by 3

rxcklesskaisher:

no c

aevans182:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @rxcklesskaisher its D \(\color{#0cbb34}{\text{End of Quote}}\) no try inputting the answer choices for x

snowflake0531:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @rxcklesskaisher no c \(\color{#0cbb34}{\text{End of Quote}}\) yeppiiiii

rxcklesskaisher:

finally lord jesus

aevans182:

yes its c